Balance Sheet (Partiall: Estimated Residual Depreciation Deprsciation 7,200 8 years Nov. 3, 2018 Dec. 1, 2018 Dec, 1, 2018 120.000 Building 6,720 8 years 70 1 600 Please fill out the balance sheet
Answer #1

Answer

  • All working forms part of the answer
  • Requirement

Property, Plant & Equipment:

Land

$             120,000

Canoes

$                13,920

Less: Accumulated Depreciation Canoes

$                   1,090

$               12,830

Building

$              169,000

Less: Accumulated Depreciation Building

$                   6,300

$             162,700

Computer

$                   9,000

Less: Accumulated Depreciation Computer

$                   1,000

$                 8,000

Office Furniture

$                   5,800

Less: Accumulated Depreciation Office Furniture

$                      140

$                 5,660

Property, Plant & Equipment, NET

$             309,190

  • Workings

---Canoes

Working

Acquired on 3 Nov 2018

Acquired on 2 Dec 2018

A

Depreciation expense per month

$                  75

$                      70

B

No. of months till 30 June 2019

8

7

C = A x B

Accumulated Depreciation - Canoes

$               600

$                    490

---Building: $900 per month
Dec 2018 to 30 June 2019 = 7 months
Accumulated Depreciation = 900 x 7 = $ 6,300

--Computer

A

Cost

$              9,000.00

B

Residual Value

$                  900.00

C=A - B

Depreciable base

$              8,100.00

D

Life [in years]

6

E=C/D

Annual SLM depreciation

$              1,350.00

F=E/C

SLM Rate

16.67%

G=F x 2

DDB Rate

33.33%

Year

Beginning Book Value

Depreciation rate

Depreciation expense

1

$                9,000.00

33.333333%

$ 1000.00 [9000 x 33.33333% x 4months /12months]

---Furniture

Accumulated depreciation = ((5800 – 1600)/10 years ) x 4months/12 months
= $ 140